### v2

# F-Bench: Rethinking Human Preference Evaluation Metrics for Benchmarking Face Generation, Customization, and Restoration

Abstract
Recent artificial intelligence (AI) generative models have demonstrated remarkable capabilities in image production, and have been widely applied to face image generation, customization, and restoration. However, many AI-generated faces (AIGFs) still suffer from issues such as unique distortions, unrealistic details, and unexpected identity shifts, underscoring the need for a comprehensive quality evaluation method for AIGFs. To this end, we introduce FaceQ, the first comprehensive AI-generated Face image database with fine-grained Quality annotations aligned with human preferences, which consists of 12K images and 491K ratings across multiple dimensions. Using the FaceQ database, we establish F-Bench, a benchmark for comparing and evaluating face generation, customization, and restoration models, highlighting strengths and weaknesses across various prompts and evaluation dimensions. Additionally, we assess the performance of existing image quality assessment (IQA) methods on FaceQ, and further propose a large multimodal model (LMM) based Face quality Evaluator (F-Eval) to accurately assess the multi-dimensional quality of generated faces in a one-for-all manner. Extensive experimental results demonstrate the state-of-the-art performance of our F-Eval. The project page is: https://mediax-sjtu.github.io/F-Bench/.
1 Introduction
| Database | Domain | Source | Images | Scores | Dimensions |
| AGIQA-3K (TCSVT2024) [51] | General | AI-generated | 2,982 | MOS | Perception, Alignment |
| AIGCIQA (CVPR2024) [89] | General | AI-generated | 2,400 | MOS | Quality, Authenticity, Correspondence |
| RichHF-18K (CVPR2024) [59] | General | AI-generated | 17,760 | MOS | Plausibility, Aesthetics, Text-image Alignment, Overall |
| PIQ23 (CVPR2023) [13] | Face | Real | 5,116 | Pair | Overall, Details, Exposure |
| FaceQ | Face | AI-generated, customized and restored | 12,255 | MOS | Quality, Authenticity, ID Fidelity, Correspondence |
Generative models, including VAE [47], GAN [26], and diffusion models [33, 82, 74], have made remarkable progress in recent years, and have been extended to the fields of text-to-image generation [3], text-driven image editing [44], image restoration [102], etc, leading to high-quality artificial intelligence generated images (AIGIs). As an important aspect of AIGIs, many studies have also investigated face generation [103], face customization [57, 111], and face restoration [62, 90], producing creative, high-quality face images. Despite these advances, many AI-generated faces (AIGFs) still suffer from perceptual deficiencies.
Many natural face image quality assessment databases have been established in the literature [48, 51, 89, 59]. However, natural images are mainly degraded by distortions such as noise, blur, compression [113, 66, 23], etc., while AIGIs generally suffer from unique distortions including unrealistic structures and text-image inconsistencies [59, 89], etc. Recently, many studies have investigated human preferences for AIGIs [59, 89, 50, 39, 106]. However, these general AIGI quality assessment works may not be well aligned with human preferences on AIGFs, since the perceptual quality of AIGFs can be significantly affected by small deviations in facial attributes,(e.g., eye size, skin texture, identity shifts), while the number of AIGFs in these AIGI databases is limited. Moreover, current AIGI quality assessment works mainly focus on the image generation, while human preferences for face customization have rarely been considered, where identity fidelity is significant.
The above limitation highlights the necessity of establishing a comprehensive quality evaluator for AIGFs. To this end, we propose FaceQ, the first AI-generated Face image Quality assessment database with three popular tasks. Specifically, FaceQ contains 12K+ images obtained from 29 generative models across three tasks: 1) face generation, 2) face customization, and 3) face restoration, as shown in Fig.1. Based on the generated images, we collect 491K+ quality ratings from multiple dimensions including quality, authenticity, identity (ID) fidelity, and text-image correspondence, respectively, finally forming 32K+ mean opinion scores (MOSs). Using FaceQ, we establish a human preference benchmark, i.e., F-Bench, for face generation, customization, and restoration models, and thoroughly analyze the generation performance across different evaluation perspectives, models, and categories. In addition, we also compare the correlation of existing image quality assessment (IQA), face quality assessment (FQA), AIGI quality assessment, and preference evaluation metrics with human preferences.
Furthermore, we present F-Eval, the first generated Face image quality Evaluation method designed to perform comprehensive multi-dimensional quality evaluation for generated, customized, and restorted faces in a one-for-all manner. Specifically, F-Eval leverages large multimodal models (LMMs) [11] and adopts instruction tuning [63] for preference-related vision-language adaption. We also apply low-rank adaptation (LoRA) [35] techniques for more efficient feature refining. Extensive experimental results demonstrate that F-Eval outperforms existing quality assessment methods in terms of multiple dimensions relevant to human preference for all three tasks. The main contributions of our work are highlighted as follows:
We construct FaceQ, the first generated face quality assessment database consisting 491K+ human preference ratings from quality, authenticity, ID fidelity, correspondence dimensions, on 12K+ generative face images across three tasks.
We establish F-Bench, a comprehensive human preference benchmark for 29 face generation, customization, and restoration methods. It also extensively evaluates the alignment of 26 existing quality assessment methods with human preferences.
We present F-Eval, a LMM-based quality assessment model to predict multi-dimensional human preference on generated faces. Extensive experiments demonstrate the state-of-art QA performance of the proposed F-Eval.
2 Related Work
Face Generation & Customization & Restoration.
(1) For face generation models, earlier works such as GANs [26, 43, 87] have demonstrated significant influences. Recently, diffusion models [74, 77, 73] have achieved rapid developments in text-to-image generation. These models can be categorized into pixel-space diffusion models [72, 73, 76, 1], and latent space diffusion models [74, 58, 71, 17]. (2) For face customization, this is a recently defined task that aims to create images of a specific person with the input of an identity reference image and prompts [101]. It can be divided into two categories including conventional test-time optimization [75, 24, 49] and tuning-free customization methods utilizing pre-trained diffusion models [104, 111, 57, 20, 30, 94].
(3) For face restoration, which aims to restore high-quality face images from their low-quality counterparts. Some traditional restoration methods utilize geometric priors [19, 46, 16] or reference priors [54, 55] while in recent years, exploring generative prior such as GAN priors [97, 110, 15, 28, 116] and diffusion priors [100, 90, 62, 112, 86, 36, 37] has become a prevalent trend. Despite these improvements, the generated, customized, and restored images may suffer from various quality issues, highlighting the necessity of evaluators.
Face Image Quality Assessment Database.
Existing face image quality assessment (FIQA) can be divided into Biometric FIQA [12, 31, 60, 69] and Generic FIQA [95, 84, 13, 18].
BFIQA methods are most developed to evaluate the biometric utility for face recognization. GFIQA methods [84] focus on perceptual degradation in face restoration task. Both of these databases consist of natural face images, which has distinctly different nature with generated faces.
AIGC Quality Assessment Database.
In recent years, plenty of researchers have raised interest in assessing AI-generated images(AIGIs) quality assessment [89, 48, 51, 23, 107, 40] by building various databases as shown in Tab 1. All of these existing database contain various categories of general AIGI images, leaving a notable gap in datasets dedicated to AI-generated face images.
3 FaceQ Database Construction
In this section, we introduce the proposed generative face quality assessment database FaceQ. The database includes 12,255 images and 491,130 human opinion ratings. Our FaceQ demonstrates the merits of comprehensive generated visual content and multi-dimensional labeling, which can significantly facilitate the advancement of AIGC research.
3.1 Data Collection
3.1.1 FaceQ-Gen Subset Collection
Prompt Sources.
Models Collection.
We collect 14 representative face generation models, including Deep Floyd [1], Stable Diffusion V1.5 [74], Stable Diffusion V2.1 [74], SD-XL [71], Realistic Vision [8], Stable Cascade [70], ProtoVision V6.6 [7], SD3 [9], Playground V2.5 [52], Kolors [6], PixArt-alpha [17], Hunyuan [58], and Flux-dev [3]. Negative prompts and safe sensors are enabled. In total, FaceQ-Gen includes 4,032 generated face images.
3.1.2 FaceQ-Cus Subset Collection
Identity and Prompt Sources.
Face customization relies on a reference identity image and a guiding prompt. The reference images in FaceQ-Cus are partially self-collected, with the rest sourced from [57]. The prompts are selected and filtered from the FaceQ-Gen prompt list.
Models Collection.
3.1.3 FaceQ-Res Subset Collection
Low Quality Images Source.
(1) Synthetic: We construct two degraded pipelines following previous works [15, 98, 100, 96] to simulate real-world degradations on both in-domain data from VFHQ [93], CelebRef [56] and out-of-domain self-collected high-quality Asian face images (2) Real-world: Real-world low-quality images are sourced from various public datasets, including AgeAB [68], CelebChild [28], Wider [108], LFW [38] and MegaFace [45], ensuring the degradation diversity.
Models Collection and Face Image Restoration.
We evaluate both diffusion-prior-based methods such as DiffBIR [62], DifFace [112], StableSR [90], and GAN or Transformer-based methods GFPGAN [97], GPEN [110], CodeFormer [116], VQFR [28] and Face-SPARNet [15]. Each method is benchmarked on both the F-Res synthetic dataset and the real-world dataset. In total,FaceQ-Res include 4023 restored images, with 2007 images in the real-world case and 2016 images in the synthetic case.
3.2 Human Preference Dimension Design
The evaluation dimensions of generated faces need specific design to adequately reflect human preference, given the complexity of perceptual factors involved. To the end, we decompose “quality” into four core dimensions: Quality, Authenticity, ID Fidelity, and Correspondence.
Quality.
This dimension evaluates the overall perceptual quality of the image, considering factors such as color accuracy, blur, noise, and artifacts. It is used in all tasks.
Authenticity.
This dimension evaluates how closely the generated image resembles a natural, real-life photograph, with particular emphasis on realistic skin texture, facial details, and wrinkles. It is used in face generation task.
ID Fidelity.
This dimension evaluates how well the identity of the reference image is preserved in the customized or restored image in face customization and restoration tasks.
Correspondence.
This dimension evaluates text-to-image alignment in face generation and customization tasks.
3.3 Subjective Experiment
Main Assessment.
To ensure a comprehensive, fair, and reliable subjective assessment, 180 participants with normal or corrected-to-normal eyesight are recruited for the experiment. They first take expert-led tutorials before the main assessment. In the main assessment, participants are asked to rate the image by dragging a sliding window from 0 to 5 with a two-digit decimal from different dimensions. All images are displayed on 27-inch 4K Dell monitors, randomly presented, at their original resolution under standard lighting conditions. The experiment follows the ITU-R BT.500-14 guidelines [78] for subjective evaluations.
Subjective Score Processing.
To eliminate the inter-annotator variability, outlier detection based on Kurtosis is conducted with a rejection rate of 3% [78]. As a result, each image is rated by at least 15 subjects. These valid ratings are converted and scaled into Z-scores ranging from with the following formulas:
Here, represents the original ratings provided by the -th evaluator for the -th image, denotes the total number of images that were assessed by the -th participant in this group. Finally, the mean opinion score (MOS) of the image is computed by averaging the rescaled z-scores as follows:
where denotes the number of valid subjects for image , and denotes rescaled z-scores.
4 F-Bench: Benchmarking Current Generative Face Models
In this section, we establish a human preference benchmark F-Bench for today’s state-of-the-art face generation, customization, and restoration methods, from quality, authenticity, ID fidelity, and correspondence dimensions. This benchmark analyzes the strengths and weaknesses, providing new insights and potential directions for future work.
4.1 Perspective Analysis
Fig.3 demonstrate the MOS distribution across eight dimensions in the three subsets, highlighting distinct variations between dimensions. (1) In the face generation task, the Correspondence score is the highest among the three dimensions, indicating that existing generative models perform well in aligning with prompts. The distribution of Authenticity is skewed towards lower scores, suggesting that current face generation models still fall short in generating photorealistic face images that satisfy human preferences. (2) In face customization, models perform the worst in the Quality dimension overall, demonstrating the difficulty of injecting identity information without compromising image quality. Identity Fidelity shows a bimodal distribution, reflecting the dichotomous nature of human judgment regarding whether the generated face accurately represents the same identity. (3) In face restoration, models perform better in Quality than in Identity Fidelity.
4.2 Model-wise Comparison
Face Generation. Fig.17a presents the average MOS scores across three dimensions: Quality, Authenticity, and Correspondence. Correspondence scores show a steady increase over time, indicating consistent improvement in text comprehension. Quality scores generally rise with the release year, although Hunyuan [58] and Dreamlike [2] exhibit deviations. Hunyuan [58] oversmoothes the facial details and performs worse compared to its contemporaries, while Dreamlike [2] performs the opposite. Authenticity scores exhibit the greatest variation among models, suggesting that current face generation models place insufficient emphasis on authenticity. Except for Realistic Vision [8], SD3 [9], Flux-dev [3], and Dreamlike [2], most models lag in authenticity relative to quality, highlighting a common phenomenon of high-quality but less realistic outputs in generative models. Fig.3d,e give some examples of MOS distributions of different methods, revealing a reasonable fine-grained performance difference.
Face Customization. Fig.17b is the performance radar chart for face customization models. IP-Adapter-FaceID-Plus[5] achieves the best performance across all dimensions, followed by PhotoMaker [57]. This highlights their great robustness in identity encoders. ELITE [101] scores the lowest in all dimensions. IP-Adapter-FaceID [4] produces high-quality images but lacks accuracy in Identity Fidelity. This may be because the base model [71] introduces many AI artifacts, which affect both Identity Fidelity and Quality. InstantID [94] and FastComposer [104] score well in Identity Fidelity but underperform significantly in Quality and Correspondence, indicating strong identity preservation but limited text-driven editing capabilities. Fig.3h illustrate the model-wise MOS distributions, revealing similar trends to those observed in the radar chart.
Face Restoration. Fig.17c displays the average performance comparison of nine face restoration models. DR2 [100] exhibits the weakest performance in both Quality and Identity fidelity. This may be due to the degradation removal process of DR2 [100], which results in the loss of fine details and excessive smoothing. In contrast, StableSR [92] and CodeFormer [116] achieve the best overall performance among the nine face restoration models, successfully preserving both intricate details and accurate identity representations. When comparing the performance on real-world and synthetic data, most methods struggle to achieve satisfactory results on complex real-world degradations, except for SAPRNet [15] and DiffBIR [62]. StableSR [92] and GFPGAN [97] show strong robustness, performing consistently well on both real-world and synthetic data.
4.3 Class-wise Comparison
For the face generation task, we analyze the Mean Opinion Score (MOS) distributions of Quality, Authenticity, and Correspondence across nine prompt categories, as shown in Fig.5. The violin plots highlight distinct performance patterns across these categories. The Action, Expression, and Facial Attributes categories exhibit lower MOS values in all three dimensions, with notable variability. In the Action category, limited visibility of the face due to body movements affects the generation of fine facial details, resulting in reduced scores, especially in Quality and Correspondence. The Expression category poses challenges in maintaining spatial relationships and physical constraints, leading to distortion and artifacts in expressions with exaggerated facial movements, such as laughing or crying, thereby impacting both Quality and Authenticity. For Facial Attributes, generating images with precise attributes (e.g., specific eye color, hair style) is challenging, particularly when multiple attributes are specified, which lowers the scores. In contrast, the General and Background categories achieve the highest MOS scores across all dimensions, as simpler or less specific prompts are easier for the model to interpret accurately. The Clothes and Accessories categories also perform well, indicating that the model handles clothes and accessory prompts effectively. Overall, prompts with specific or complex attribute requirements tend to result in lower MOS scores, while general or simple prompts yield consistently high scores across dimensions. This can be attributed to the absence of high-quality fine-grained text-face pair datasets during the training of generative models.
| Task | Face Generation | Face Customization | ||||||||||||||||
| Dimension | Quality | Authenticity | Correspondence | Quality | ID Fidelity | Correspondence | ||||||||||||
| Model | SRCC | KRCC | PLCC | SRCC | KRCC | PLCC | SRCC | KRCC | PLCC | SRCC | KRCC | PLCC | SRCC | KRCC | PLCC | SRCC | KRCC | PLCC |
| NIQE [67] | 0.1726 | 0.1142 | 0.2553 | 0.0869 | 0.0571 | 0.1346 | 0.0752 | 0.0507 | 0.1098 | 0.1365 | 0.0899 | 0.1436 | 0.1490 | 0.0991 | 0.1310 | 0.1186 | 0.0784 | 0.1245 |
| ILNIQE [114] | 0.1981 | 0.1319 | 0.3159 | 0.0675 | 0.0445 | 0.1505 | 0.0889 | 0.0591 | 0.1665 | 0.2032 | 0.1327 | 0.2479 | 0.2192 | 0.1452 | 0.2778 | 0.1878 | 0.1251 | 0.2275 |
| HOSA [105] | 0.2537 | 0.1708 | 0.2972 | 0.1103 | 0.0728 | 0.1446 | 0.122 | 0.0814 | 0.1495 | 0.1019 | 0.0688 | 0.0597 | 0.0753 | 0.0506 | 0.0552 | 0.1036 | 0.0696 | 0.0748 |
| BPRI-PSS [64] | 0.2362 | 0.1595 | 0.2413 | 0.1700 | 0.1142 | 0.1772 | 0.1673 | 0.1121 | 0.1647 | 0.0238 | 0.0166 | 0.0314 | 0.1425 | 0.0972 | 0.1758 | 0.1100 | 0.0729 | 0.1067 |
| FISBLIM [27] | 0.0901 | 0.0595 | 0.0433 | 0.0028 | 0.0003 | 0.0175 | 0.1000 | 0.0665 | 0.0196 | 0.1614 | 0.1024 | 0.1690 | 0.0256 | 0.0173 | 0.0126 | 0.0852 | 0.0559 | 0.0854 |
| ArcFace[22] | N/A | N/A | N/A | N/A | N/A | N/A | N/A | N/A | N/A | 0.3099 | 0.2094 | 0.3605 | 0.5062 | 0.3439 | 0.5572 | 0.1837 | 0.1243 | 0.2244 |
| SER-FIQ[85] | 0.1552 | 0.1034 | 0.1561 | 0.0470 | 0.0314 | 0.0575 | 0.1130 | 0.075 | 0.1110 | -0.1241 | -0.080 | -0.1241 | -0.0109 | -0.0078 | 0.0042 | -0.1680 | -0.1121 | -0.1911 |
| DSL-FIQA[18] | 0.5606 | 0.3927 | 0.5945 | 0.3156 | 0.2096 | 0.3960 | 0.3706 | 0.2514 | 0.4355 | 0.3615 | 0.2450 | 0.3651 | 0.2429 | 0.1652 | 0.2519 | 0.1040 | 0.0703 | 0.0974 |
| IFQA[41] | 0.2313 | 0.1549 | 0.3261 | 0.1597 | 0.1062 | 0.2097 | 0.1597 | 0.1077 | 0.2260 | 0.2401 | 0.1614 | 0.2220 | 0.1503 | 0.0973 | 0.1739 | 0.3731 | 0.2546 | 0.3941 |
| CNNIQA* [42] | 0.4219 | 0.2895 | 0.3510 | 0.3037 | 0.2054 | 0.2973 | 0.2824 | 0.1931 | 0.2595 | 0.6244 | 0.4376 | 0.6620 | 0.6511 | 0.4441 | 0.6381 | 0.3765 | 0.2545 | 0.3264 |
| VGG16* [81] | 0.5846 | 0.4141 | 0.6006 | 0.5010 | 0.3508 | 0.5181 | 0.4349 | 0.2994 | 0.4723 | 0.7217 | 0.5324 | 0.7857 | 0.6735 | 0.4717 | 0.6636 | 0.5635 | 0.3932 | 0.5589 |
| VGG19* [81] | 0.5728 | 0.4043 | 0.5644 | 0.4681 | 0.3248 | 0.4479 | 0.4171 | 0.2873 | 0.4754 | 0.7952 | 0.6031 | 0.8096 | 0.7544 | 0.5417 | 0.7555 | 0.6585 | 0.4742 | 0.6465 |
| ResNet18* [29] | 0.6150 | 0.4417 | 0.6660 | 0.4600 | 0.3160 | 0.4604 | 0.4829 | 0.3360 | 0.4628 | 0.7966 | 0.6013 | 0.8006 | 0.7473 | 0.5335 | 0.7656 | 0.6308 | 0.4488 | 0.6191 |
| ResNet34* [29] | 0.6092 | 0.4354 | 0.6537 | 0.5564 | 0.3949 | 0.5401 | 0.4410 | 0.3065 | 0.5207 | 0.8141 | 0.6211 | 0.8151 | 0.7969 | 0.5878 | 0.8157 | 0.5708 | 0.4000 | 0.5810 |
| HyperIQA* [83] | 0.6333 | 0.4545 | 0.6591 | 0.6106 | 0.4406 | 0.6042 | 0.4235 | 0.2912 | 0.4240 | 0.8524 | 0.6631 | 0.8205 | 0.8419 | 0.6463 | 0.8588 | 0.6628 | 0.4757 | 0.6454 |
| TReS* [25] | 0.7766 | 0.5867 | 0.8047 | 0.6458 | 0.4703 | 0.6579 | 0.5708 | 0.4038 | 0.6270 | 0.8883 | 0.7083 | 0.8951 | 0.8330 | 0.6409 | 0.8658 | 0.7393 | 0.5497 | 0.7400 |
| MANIQA* [109] | 0.7871 | 0.5951 | 0.8150 | 0.7308 | 0.5436 | 0.7794 | 0.6278 | 0.4443 | 0.7066 | 0.8970 | 0.7185 | 0.8952 | 0.8590 | 0.6654 | 0.8796 | 0.7849 | 0.5941 | 0.7822 |
| PromptIQA* [14] | 0.7823 | 0.5907 | 0.8245 | 0.7232 | 0.5327 | 0.7558 | 0.6514 | 0.4698 | 0.7065 | 0.9071 | 0.7315 | 0.8821 | 0.8693 | 0.6792 | 0.8876 | 0.8348 | 0.6342 | 0.8281 |
| CLIPScore [32] | 0.0967 | 0.0642 | 0.1315 | 0.0640 | 0.0408 | 0.0817 | 0.2396 | 0.1591 | 0.2733 | 0.3126 | 0.2106 | 0.3185 | 0.1816 | 0.1218 | 0.1811 | 0.7008 | 0.5098 | 0.6986 |
| BLIPScore [53] | 0.1306 | 0.0872 | 0.1929 | 0.0935 | 0.0613 | 0.1569 | 0.2051 | 0.1364 | 0.2963 | 0.3073 | 0.2063 | 0.3136 | 0.1213 | 0.0796 | 0.1266 | 0.6707 | 0.4797 | 0.6123 |
| ImageReward [106] | 0.3849 | 0.2631 | 0.4358 | 0.2606 | 0.1736 | 0.3181 | 0.5155 | 0.3542 | 0.5871 | 0.3232 | 0.2166 | 0.3196 | 0.1905 | 0.1270 | 0.1839 | 0.7714 | 0.5673 | 0.7740 |
| MINTIQA[91] | 0.7002 | 0.5160 | 0.8047 | 0.6217 | 0.4450 | 0.6826 | 0.5491 | 0.3891 | 0.6584 | 0.5378 | 0.3775 | 0.5574 | 0.4088 | 0.2816 | 0.4041 | 0.2886 | 0.1932 | 0.2959 |
| MINTIQA* [91] | 0.8312 | 0.6474 | 0.8974 | 0.8177 | 0.6306 | 0.8511 | 0.7908 | 0.5991 | 0.8667 | 0.8904 | 0.7092 | 0.8816 | 0.8524 | 0.6697 | 0.8755 | 0.8391 | 0.6413 | 0.8318 |
| F-Eval (Ours) | 0.8486 | 0.6670 | 0.9085 | 0.8312 | 0.6585 | 0.8578 | 0.8471 | 0.6637 | 0.9106 | 0.9462 | 0.7961 | 0.9461 | 0.9188 | 0.7640 | 0.9322 | 0.9460 | 0.7959 | 0.9457 |
5 F-Eval Method
In this section, we introduce our one-for-all generated face quality assessment method, F-Eval, towards predicting quality scores from multiple dimensions for three tasks using one model.
5.1 Overall Architecture
The overall architecture of F-Eval is depicted in Fig. 6. As shown on the left, F-Eval can assess four dimensions: quality, authenticity, text-image correspondence, and identity fidelity, with multimodal inputs, i.e.text prompts, image and optional reference image, to output quality-related answers. On the right, the input image(s) can originate from all three tasks. F-Eval employs a vision encoder and a face encoder to extract visual features, with projectors projecting both into a unified language space. User prompts are encoded into textual embeddings through a tokenizer. These multimodal features are concatenated and fed into a pre-trained large language model for the final output.
5.2 Model Design
Vision Encoding.
For input image(s), F-Eval adopt a vision encoder to extract global features and a face encoder captures face features related to facial structure and identity. Specifically, the vision encoder, Qwen-VL-2.5-ViT [11], is tuned on FaceQ to adjust different dimensions. Face encoder is Arcface [22], which is frozen. Two two-layer multi-layer perception projectors then map the features .
Feature Fusion via the LLM.
For input prompts, it is tokenized into text tokens and then concatenated with visual tokens. Here, Qwen-VL2.5-7B [11] is employed as the pre-trained LLM backbone to combine visual tokens and text tokens to perform multi-modal learning.. Finally, the output features of the LLM are decoded with a text decoder and then projected to the prompt-corresponded quality space.
5.3 Multi-Expert LoRA Tuning
Instruction Tuning.
Previous IQA methods are often designed for specific dimensions [109, 115], addressing only one aspect of the problem or requiring retraining of model weights for different dimensions. Leveraging the generalization ability of LLMs, instruction tuning allows for the integration of different dimensions within a unified framework. Therefore, we construct diverse instruction data for the four facial quality evaluation dimensions in the Face-Q dataset and fine-tune F-Eval by adjusting the visual encoder, projector, and LLM parameters to adapt it to the specific task. As a result, F-Eval is capable of evaluating facial quality using unified weights.
Mixture of LoRA Experts (MoLE).
Tuning a holistic quality model for different dimensions using a single LoRA expert introduces inherent interference between dimensions. Hence, F-Eval integrates four LoRA experts for four dimensions when tuning vision encoder and LLM, constituting MoLE modules. A router is trained to classify input prompts into the four dimensions and output the corresponding dimension ID . Specifically, for a frozen linear layer with input and weight matrix , a LoRA expert is trained to learn a low-rank decomposed update:
where and are the low-rank matrices, is the rank, and is scaling factor. For prompt of certain dimension, the corresponding expert is activated based on the dimension ID , which can be formulated as:
where is the k-th LoRA expert.
6 Experiments
6.1 Implementation Details
The LLM used in F-Eval is Qwen-2.5VL [11], with an input token channel dimension of 3584. The vision encoder, face encoder, and LLM are all frozen during training. The ranks of LoRA experts are 16,32,16,16, respectively. The LoRA alphas are set equal to LoRA ranks. LoRA dropout is set as 0.05. The model is trained on 2 NVIDIA GeForce RTX 3090 GPUs with flash-attn [21] enabled. We use AdamW Optimizer to train the network for 10 epochs with a batch size of 4 and learning rate .
6.2 Comparison with existing QA Models
We conduct comprehensive experiments on the proposed FaceQ database to evaluate the effectiveness of both current quality assessment methods and our proposed methods, as shown in Tab.2, Tab.3, and Tab.4. The evaluation metrics are Spearman rank-order correlation coefficient (SRCC), Kendall rank order correlation coefficient (KRCC), and Pearson linear correlation coefficient (PLCC), which assess the correlation between predicted scores by the QA methods and the MOS scores provided by humans. Higher metrics indicated better alignment with human preferences. The comparison QA models falls into four categories: 1) Traditional IQA Models. NR-IQA models include NIQE [67], ILNIQE [114], HOSA [105], BPRI-PSS [64], FISBLIM [27]. FR-IQA models include LPIPS [115], SSIM [99] and PSNR [34]. 2) Classical Deep learning-based IQA Models. This category includes CNNIQA [42], VGG16 [81], VGG19 [81], ResNet18 [29], ResNet34 [29], HyperIQA [83], TReS [25], MANIQA [109],PromptIQA[14]. All models are retrained on FaceQ with an 80:20 training/testing split. 3) AIGC IQA Models. This category includes pre-trained vision-language models, including CLIPScore [32], BLIPScore [53], zero-shot metric ImageReward [106], and the AIGC-specific QA model MINTIQA [91]. 4) Face IQA Models. This category includes SER-FIQ [85], DSL-FIQA [18], ArcFace [22], IFQA[41].
6.3 Performance Analysis
| Dimension | Quality | ID Fidelity | ||||
| Model | SRCC | KRCC | PLCC | SRCC | KRCC | PLCC |
| BMPRI[65] | 0.2161 | 0.1428 | 0.1862 | 0.1113 | 0.0739 | 0.1069 |
| NIQE[67] | 0.1798 | 0.1213 | 0.1642 | 0.2193 | 0.1480 | 0.2118 |
| ILNIQE[114] | 0.2505 | 0.1727 | 0.3180 | 0.2696 | 0.1857 | 0.3321 |
| HOSA[105] | 0.3758 | 0.2522 | 0.3748 | 0.3418 | 0.3455 | 0.2301 |
| BPRI-LSSn [64] | 0.1398 | 0.0957 | 0.0999 | 0.1244 | 0.0837 | 0.0973 |
| BPRI [64] | 0.1637 | 0.1065 | 0.1480 | 0.0195 | 0.0121 | 0.0281 |
| FISBLIM [27] | 0.1589 | 0.1089 | 0.2038 | 0.1054 | 0.0709 | 0.1715 |
| LPIPS[115] | 0.3909 | 0.2658 | 0.4643 | 0.4372 | 0.3018 | 0.4870 |
| SSIM[99] | 0.1017 | 0.0675 | 0.0488 | 0.0933 | 0.0601 | 0.0504 |
| PSNR [34] | 0.0198 | 0.0113 | 0.0536 | 0.0986 | 0.0662 | 0.1212 |
| ArcFace[22] | 0.3466 | 0.2327 | 0.3774 | 0.5728 | 0.4030 | 0.5762 |
| SER-FIQ [25] | 0.3813 | 0.2584 | 0.3746 | 0.3951 | 0.2641 | 0.3707 |
| DSL-FIQA [18] | 0.6387 | 0.4508 | 0.6528 | 0.4882 | 0.3360 | 0.5379 |
| IFQA [41] | 0.2273 | 0.1527 | 0.3348 | 0.3214 | 0.2185 | 0.3874 |
| CNNIQA* [42] | 0.4661 | 0.3197 | 0.3836 | 0.4678 | 0.3216 | 0.4288 |
| ResNet18* [29] | 0.7254 | 0.5295 | 0.6982 | 0.6662 | 0.4802 | 0.6527 |
| ResNet34*[29] | 0.6693 | 0.4916 | 0.6550 | 0.5905 | 0.4170 | 0.6084 |
| VGG16* [81] | 0.7487 | 0.5513 | 0.7101 | 0.6177 | 0.4413 | 0.6381 |
| VGG19* [81] | 0.6994 | 0.5126 | 0.6722 | 0.6038 | 0.4278 | 0.5999 |
| HyperIQA* [83] | 0.8269 | 0.6323 | 0.8228 | 0.7615 | 0.5585 | 0.7471 |
| MANIQA* [109] | 0.8287 | 0.6403 | 0.8778 | 0.7704 | 0.5660 | 0.7619 |
| TReS* [25] | 0.8656 | 0.6830 | 0.8622 | 0.7951 | 0.5879 | 0.7829 |
| PromptIQA* [14] | 0.8519 | 0.6661 | 0.8733 | 0.8220 | 0.6275 | 0.8254 |
| F-Eval (Ours) | 0.8692 | 0.6855 | 0.9009 | 0.8507 | 0.6731 | 0.8726 |
| Dimension | Quality | ID Fidelity | ||||
| Model | SRCC | KRCC | PLCC | SRCC | KRCC | PLCC |
| BMPRI[65] | 0.1582 | 0.1093 | 0.2799 | 0.0149 | 0.0108 | 0.1386 |
| NIQE[67] | 0.1783 | 0.1186 | 0.1962 | 0.1245 | 0.0810 | 0.1701 |
| ILNIQE[114] | 0.2230 | 0.1610 | 0.2291 | 0.1964 | 0.1362 | 0.2340 |
| HOSA[105] | 0.3927 | 0.4277 | 0.2654 | 0.3328 | 0.2234 | 0.3813 |
| BPRI-LSSn [64] | 0.2715 | 0.1840 | 0.3022 | 0.2059 | 0.1380 | 0.2475 |
| BPRI [64] | 0.1920 | 0.1313 | 0.2200 | 0.0846 | 0.0570 | 0.1136 |
| FISBLIM [27] | 0.2512 | 0.1705 | 0.2725 | 0.2341 | 0.1581 | 0.2699 |
| ArcFace[22] | 0.2371 | 0.1565 | 0.4183 | 0.4956 | 0.3470 | 0.6330 |
| SER-FIQ [25] | -0.0482 | -0.0299 | 0.1023 | 0.0421 | 0.0299 | 0.1584 |
| DSL-FIQA [18] | 0.6509 | 0.4641 | 0.6494 | 0.4840 | 0.3362 | 0.5118 |
| IFQA [41] | 0.0251 | 0.0210 | 0.1221 | 0.0893 | 0.0603 | 0.2086 |
| CNNIQA* [42] | 0.3650 | 0.2506 | 0.4705 | 0.3252 | 0.2197 | 0.3112 |
| VGG16*[81] | 0.4996 | 0.3439 | 0.4269 | 0.4117 | 0.2859 | 0.4622 |
| VGG19* [81] | 0.5871 | 0.4181 | 0.5133 | 0.3054 | 0.2044 | 0.3594 |
| ResNet18* [29] | 0.5977 | 0.4221 | 0.5722 | 0.5554 | 0.3909 | 0.6161 |
| ResNet34* [29] | 0.6259 | 0.4475 | 0.6168 | 0.5036 | 0.3539 | 0.6136 |
| HyperIQA* [83] | 0.7926 | 0.5965 | 0.7926 | 0.7783 | 0.5845 | 0.8163 |
| MANIQA* [109] | 0.8356 | 0.6482 | 0.8467 | 0.7597 | 0.5655 | 0.8081 |
| TReS* [25] | 0.8287 | 0.6391 | 0.8501 | 0.7287 | 0.5441 | 0.8009 |
| PromptIQA* [14] | 0.8115 | 0.6186 | 0.8645 | 0.7398 | 0.5526 | 0.8297 |
| F-Eval(Ours) | 0.8448 | 0.6577 | 0.8705 | 0.7957 | 0.6057 | 0.8366 |
Traditional IQA Models.
Traditional IQA methods, including both NR-IQA and FR-IQA models, generally perform poorly across all tasks and dimensions, with low SRCC, KRCC, and PLCC scores.
Classical Deep Learning-based IQA Models.
AIGC IQA Models.
CLIPScore [32] BLIPScore [53], and ImageReward [106] exhibit weak correlations with human preference, likely due to their focus on high-level semantics rather than low-level details crucial for image quality assessment. MINTIQA [91] achieves best performance among current AIGC IQA methods, but its performance is still unsatisfactory, particularly on the customization task. Fine-tuning on our dataset leads to significant improvements, highlighting the great potential of FaceQ.
Face IQA Models.
ArcFace similarity [22] demonstrates its ability to extract identity-related features. However, it lacks perceptual sensitivity for image quality, leading to poor performance in quality assessment. DSL-FIQA [18], as the latest method, achieves the best performance in predicting quality scores among face IQA methods, but it is still far from satisfactory. SER-FIQ fails to align with human perception across most dimensions and tasks. Experiments confirm that these natural face evaluation algorithms struggle to assess the complex characteristics of AIGFs, highlighting the need for an effective AIGF evaluator.
F-Eval (Ours).
Our method achieves optimal performance across all tasks and dimensions, with the most significant gain in the customization task, as existing methods rarely focus human preferences in face customization. Among the four dimensions, the improvements in correspondence and ID fidelity stand out, thanks to dual-stream vision encoding and the powerful LLM backbone for text understanding and modality fusion. The MoLE technique allows the model to learn distinct features for each dimension.
6.4 Ablation Study
We conducted the ablation study to validate the utility of the core components of F-Eval in Tab. 5.
Effectiveness of Vision Encoding.
1) The effectiveness of the face encoder is demonstrated by comparing the second and last rows. Without the face encoder, a significant performance drop is observed, indicating that the extracted features help the LLM interpret face images. 2) Removing the projector leads to a dramatic performance drop.
Effectiveness of MoLE.
We conducted three sets of LoRA adaptation experiments: fine-tuning a single LoRA for all dimensions and tasks, fine-tuning three LoRA experts for three tasks, and fine-tuning four LoRA experts for four dimensions. The results show that fine-tuning multiple dimensions with a single LoRA leads to performance deterioration, which further confirms the independence of the characteristics across different dimensions.
| Feature & Strategy | Face Generation | Face Customization | Face Restoration syn | Face Restoration rw | |||||||||||
| Vision | Projector | Face | Single LoRA | Task LoRA | Dim. LoRA | Quality | Authen. | Corres. | Quality | ID. | Corres. | Quality | ID. | Quality | ID. |
| ✔ | ✔ | ✔ | 0.808 | 0.750 | 0.846 | 0.924 | 0.843 | 0.927 | 0.777 | 0.776 | 0.707 | 0.478 | |||
| ✔ | ✔ | ✔ | 0.815 | 0.752 | 0.835 | 0.924 | 0.855 | 0.921 | 0.812 | 0.783 | 0.716 | 0.603 | |||
| ✔ | ✔ | ✔ | ✔ | 0.793 | 0.776 | 0.819 | 0.941 | 0.885 | 0.942 | 0.838 | 0.823 | 0.802 | 0.671 | ||
| ✔ | ✔ | ✔ | ✔ | 0.823 | 0.780 | 0.838 | 0.942 | 0.873 | 0.942 | 0.824 | 0.826 | 0.754 | 0.638 | ||
| ✔ | ✔ | ✔ | ✔ | 0.849 | 0.831 | 0.847 | 0.946 | 0.919 | 0.946 | 0.869 | 0.851 | 0.845 | 0.796 | ||
7 Conclusion
In this work, we conduct a pioneering generative face image quality assessment study. Specifically, we introduce FaceQ, the first generated face quality assessment database comprising 12,255 face images and 491,130 human preference ratings. Then, we establish F-Bench, a benchmark for evaluating the capabilities of existing face generation, customization, and restoration models, while also reveals the limitations of current Image Quality Assessment (IQA) and Face Quality Assessment (FQA) methods. Furthermore, we propose an evaluation model F-Eval to accurately predict human preferences on generated face images, which demonstrates state-of-the-art performance. We believe this work will contribute to advancing generative models and help overcome current limitations in AIGC community.
Acknowledgment
This work is supported by the National Natural Science Foundation of China under Grants 62271308, 62401365, 62225112, 62132006, U24A20220, STCSM under Grants 24ZR1432000, 24511106902, 24511106900, 22511105700, 22DZ2229005, 111 plan under Grant Number BP0719010, China Postdoctoral Science Foundation under Grants BX20250411, 2025M773473, and State Key Laboratory of UHD Video and Audio Production and Presentation.
Supplementary Material
Limitations and Social Impact.
First, we are mindful of privacy concerns related to face datasets, and all data collection and sharing adhere to relevant privacy policies. Second, while the maximum resolution in our dataset is limited to 1024 pixels, we recognize that future work may incorporate higher-resolution images as generative models continue to advance.
8 FaceQ: Dataset Construction
8.1 Face Generation, Customization and Restoration Models
Model Implementation Details.
Tab. 8.1 provides a comprehensive summary of models evaluated for face generation, editing, and restoration, including the model links, released dates, the resolution, and the backbone architectures. (1) Face generation. All the generation models are inference by pre-trained checkpoints in their default resolutions and hyper-parameters. Specifically, Stable Diffusion V1.5 [74], DreamLike [2], and RealisticVision [8] support high-resolution generation, such as 1024 1024, but we utilize their default training resolution due to severe subject repetition phenomenon. Deep Floyd [1] is a 3-stage pixel space diffusion model, here we only consider the third-stage results. For the dynamic step sampling, the number of steps per stage for StableCascade [70] and Deep Floyd [1] is reduced to the quarter. Each model’s negative prompt is configured to exclude “anime” and “semi-realistic” outputs by using terms like “worst quality”, “low quality”, “illustration”, “3D”, “2D”, “painting”, “cartoons”, “sketch”, “anime”, “animation”, “cartoon”, and “semi-realistic”. Safe sensors are enabled to filter out NSFW content. (2) Face customization. All the customization models are inference by pre-trained checkpoints in their default resolutions and hyper-parameters. FastComposer [104], originally designed for multi-subject customization, is assessed here with a single reference image as input. We use the IP-Adapter release version, including IP-Adapter-FaceID-SDXL [4] and IP-Adapter-FaceID-PlusV2 [5] (referred to as IP-Adapter-FaceID and IP-Adapter-FaceID-Plus). Their backbones are SDXL and SD-v1.5, respectively. (3) Face restoration. For DR2 [100], we follow the hyperparameter settings recommended in the original paper: for real-world inputs and for synthetic inputs.
| Category | Model | Year | Resol. | Backbone |
| Generation | Stable Diffusion V1.5 [rombach2022highresolutionimagesynthesislatent] | 2022.04 | Latent Diffusion | |
| Stable Diffusion V2.1 [rombach2022highresolutionimagesynthesislatent] | 2022.12 | Latent Diffusion | ||
| DreamLike V2.0 [2] | 2023.01 | Latent Diffusion | ||
| Deep Floyd [1] | 2023.04 | Pixel Diffusion | ||
| SD-XL [71] | 2023.06 | Latent Diffusion | ||
| PixArt-alpha [17] | 2023.11 | Latent Diffusion (DiT) | ||
| Realistic Vision V5.1 [8] | 2023.12 | Latent Diffusion | ||
| Stable Cascade [70] | 2024.02 | Latent Diffusion | ||
| Playground V2.5 [52] | 2024.02 | Latent Diffusion | ||
| ProtoVision V6.6 [7] | 2024.03 | Latent Diffusion | ||
| Hunyuan [58] | 2024.05 | Latent Diffusion (DiT) | ||
| SD3 [9] | 2024.07 | Latent Diffusion (DiT) | ||
| Kolors [6] | 2024.07 | Latent Diffusion | ||
| Flux-dev [3] | 2024.08 | Latent Diffusion (DiT) | ||
| Customization | ELITE [101] | 2023.02 | Latent Diffusion | |
| FastComposer [104] | 2023.05 | Latent Diffusion | ||
| IP-Adapter-FaceID [4] | 2023.12 | Latent Diffusion | ||
| InstantID [94] | 2023.12 | Latent Diffusion | ||
| IP-Adapter-FaceID-Plus [5] | 2023.12 | Latent Diffusion | ||
| PhotoMaker [57] | 2023.12 | Latent Diffusion | ||
| Restoration | SPARNet [15] | 2020.12 | GAN | |
| GPEN [110] | 2021.05 | GAN | ||
| GFPGAN [97] | 2021.06 | GAN | ||
| CodeFormer [116] | 2022.08 | VQ | ||
| VQFR [28] | 2022.07 | VQ | ||
| DifFace [112] | 2022.12 | Pixel Diffusion | ||
| DR2 [100] | 2023.05 | Pixel Diffusion | ||
| DiffBIR [62] | 2023.08 | Latent Diffusion | ||
| StableSR [92] | 2024.06 | Latent Diffusion |
Prompt Examples.
| Category | Prompt Examples |
| General | a photo of a woman |
| a photo of a middle-eastern man | |
| Clothing | a woman wearing a purple wizard outfit |
| a man wearing a hoodie with green stripes | |
| Accessory | a man with black hair styled in a top bun |
| an old woman with a vintage hairpin | |
| Action | a woman coding in front of a computer |
| a man playing the violin | |
| Expression | a man crying disappointedly, with tears flowing |
| a woman looking shocked, mouth wide open | |
| Background | a woman laughing on the lawn |
| a young woman with a colorful umbrella stands near a crowd | |
| View | a man wearing a doctoral cap, upper body, with the left side of the face facing the camera |
| a man playing the guitar in the view of left side | |
| Style | instagram photo, portrait photo of a man, perfect face, natural skin, hard shadows, film grain |
| editorial portrait of a man posing dramatically, sharp lighting, fashion magazine style | |
| Facial Attributes | a young girl with large round blue eyes, a flat nose bridge, and purple lipstick |
| A man with narrow black eyes, a high nose bridge, a thick beard, and fair skin |
The face-centric prompts used for face generation can be categorized into nine classes. Tab. 7 presents two example prompts for each category due to space constraints. We ensured equal numbers of prompts for male and female subjects. The prompts for face customization are a subset of those used for face generation. Degradation Scheme. We construct two synthetic degradation pipelines to mimic real-world degradation. The first is first order pipeline following previous works [15, 98, 100] which can be expressed as
High-quality images are degraded through a series of operations, including blurring, downsampling, additive Gaussian noise, and JPEG compression, with respective probabilities of 70%, 100%, 20%, and 70%. The blur kernel is randomly selected from Gaussian, Average, Median, and Motion blur. The interpolation method is randomly selected from Nearest, Linear, Area, and Cubic interpolation. The downsampling scale factor is randomly chosen from 4, 8, or 16. The second is a second-order degradation pipeline from previous work [96].
Blur, resizing, noise, and JPEG compression are conducted in several orders, along with a sinc filter to simulate common ringing and overshoot artifacts. We used these two pipelines to generate 50% and 50% of the synthetic low-quality images, respectively.
8.2 Additional Examples of FaceQ Database
Figure 7, Figure 8, and Figure 9 present additional examples from the FaceQ-Gen, FaceQ-Cus, and FaceQ-Res subsets, respectively. Each row corresponds to a specific generative model, showcasing the extensive diversity of content covered by the FaceQ dataset.
8.3 Quantitative Analysis of FaceQ Database
We selected four low-level features—brightness, contrast, colorfulness, and sharpness—to quantitatively assess the content diversity of the FaceQ database. Fig. 10 illustrates the kernel distribution curves for each selected feature across the three subsets. The results indicate that the images in each subset exhibit a wide range of contrast, colorfulness, and sharpness. However, the FaceQ-Cus subset demonstrates a narrower distribution in terms of brightness compared to the other subsets.
We further calculate the relative range and coverage uniformity of the three subsets across these selected features. The relative range is defined as:
where denotes the distribution of dataset on feature. refers to the maximum value of feature across all datasets. The coverage uniformity is calculated as the entropy of the B-bin histogram of for each subset, using the following formula:
where denotes the normalized number in bin at feature for dataset.
Fig. 11 presents a quantitative comparison of uniformity and relative range. A higher coverage uniformity indicates a more uniform feature distribution within the database, while a higher relative range reflects greater intra- and inter-dataset differences. It can be observed that all three subsets exhibit a diverse range and a uniform distribution across the four low-level features.
9 FaceQ: Subjective Experiments
9.1 Implementation Details
Fig. 12 presents screenshots of the user rating interfaces for the four tasks. In the generation task, as shown in Figure 12 (a), participants are asked to rate images on a scale of 0 to 5 based on quality, authenticity, and correspondence. Prompts are displayed beneath the candidate images, accompanied by translations into the participants’ native languages. In the customization task, as shown in Fig. 12 (b), the reference image is displayed on the left, with prompts and translation shown below. In Fig. 12 (c), the candidate image appears on the left, while the corresponding low-quality reference image is on the right. In Figure 12 (d), both the low-quality image and the ground truth are displayed in synthetic scenarios. Each subset in FaceQ was randomly divided into four groups, each containing approximately 1,000 images. Participants were compensated $14 for completing each group of experiments according to [80]. At last, 3% invalid data are removed and no subject is removed.
9.2 Subjective Evaluation Examples
Fig. 13 provides a visual supplement to the 3D scatter plots described in the main submission. Fig. 13 (a) presents the 3D scatter plot for the FaceQ-Gen subset, showcasing five representative edge points. These images, ranked from top to bottom, correspond to overall good, low correspondence, low authenticity, low quality, and overall bad. As illustrated, the MOS scores effectively and intuitively capture the strengths and weaknesses of the images, accurately reflecting human preferences across different dimensions. Similarly, Fig. 13 (b) depicts the 3D scatter plot for the FaceQ-Cus subset, highlighting another set of five representative points. These images, ranked from top to bottom, correspond to overall good, low correspondence, low identity correspondence, low quality, and overall bad. The MOS score demonstrates a significant decline in dimensions where the image exhibits poor performance. This observation further substantiates the reliability and validity of human scoring in reflecting image quality across multiple dimensions.
10 F-Bench: More Analysis
10.1 MOS Distribution
Fig. 14 illustrates the MOS distributions of all fourteen face generation models across the dimensions of quality, authenticity, and correspondence, for both full-step and 1/4-step performances. The full-step distribution plots provide a comprehensive view of the performance distribution for different methods across the three dimensions, enabling a detailed evaluation of each method’s effectiveness. In the 1/4-step distribution plots, it can be observed that models such as Stable Cascade [70], SDXL [71], and Pixart-alpha [17] exhibit high sensitivity to the reduction in step size. In contrast, models such as Flux[3] and RealisticVision[8] demonstrate relatively stable performance with minimal degradation when reducing the steps. Fig. 15 illustrates the MOS distributions for six face customization models across three dimensions: quality, identity fidelity, and correspondence. Significant variations can be observed among different models and dimensions, highlighting distinct performance characteristics. Fig. 16 illustrates the MOS distributions for all nine face restoration models across the dimensions of quality and identity fidelity, evaluated for both real-world and synthetic cases. Most models exhibit varying performance between real-world and synthetic inputs, resulting in noticeable differences in their distributions, as exemplified by SPARNet [15].
10.2 Perspectiive Analysis
To provide a clearer comparison of the strengths and weaknesses of different methods, we present the average MOS scores across various dimensions in Figure 17. For the three dimensions of face generation, authenticity exhibits the largest disparity between methods, while correspondence and quality tend to cluster around higher scores. In the face customization task, the methods show inconsistent performance in correspondence, whereas quality remains relatively balanced. For face restoration, quality-synthetic emerges as the easiest metric to achieve high scores, followed by quality-real-world. Figure 18 displays the rankings of the various methods. For face generation, Flux [3] achieves the highest performance across all three dimensions. When considering authenticity, RealisticVision [8] and SD3 [9] outperform other methods. Playground [52] ranks second only to Flux [3] in terms of correspondence, while Kolors [6] and SD3 [9] follow Flux [3] in quality. On the other hand, SDv1.5 [rombach2022high] performs the worst across all dimensions. For face customization, IP-Adapter-FaceID-Plus [5], InstantID [94], and PhotoMaker [79] excel at preserving identity information. For face restoration, CodeFormer [116] demonstrates the best performance in synthetic scenarios, while StableSR achieves the highest scores in real-world scenarios.
10.3 Class-wise Comparison
Age.
Figure 19 presents the multi-dimensional MOS distributions across three age groups (Young, Middle-aged, and Old) for face generation, face customization, and face restoration tasks. In the face generation task, the performance across age groups is relatively consistent across all dimensions. For face customization, more pronounced differences are observed, particularly in the quality scores, where older individuals exhibit larger variability. In the face restoration task, quality scores for old individuals are notably higher compared to middle-aged and young groups, while identity fidelity remains relatively consistent. These results highlight that face generation models are less sensitive to age-related factors, whereas face customization and restoration models demonstrate noticeable performance disparities among age groups, especially in dimensions such as ID Fidelity and Quality. The age and gender of the images are labeled InsightFace.
Gender.
We visualize the distribution of MOS scores in three dimensions for men and women in each dimension in Fig. 20. It can be found that the male and female categories in face generation and face customization perform consistently across all evaluation dimensions, with minimal variability observed. However, when it comes to face restoration tasks, the quality and identity fidelity of the male class are better. This suggests that generation and customization models trained on extensive datasets exhibit less gender bias than restoration models trained on smaller datasets.
11 QA Methods Implementation Details
11.1 Evaluation Metrics
We adopt three widely used metrics in IQA [88, 89]: Spearman rank-order correlation coefficient (SRCC), Pearson linear correlation coefficient (PLCC), and Kendall rank correlation coefficient (KLCC) to evaluate the performance of quality assessment methods.
SRCC., which ranges from -1 to 1, evaluates the monotonic relationship between two variables. For images, it is computed as:
Here, represents the rank of the ground truth value , while corresponds to the rank of the predicted value . When the SRCC value is higher, it signifies a stronger monotonic agreement between the ground truth and the predicted scores. PLCC. quantifies the linear correlation between predicted scores and ground truth scores and is formulated as:
where and denote the mean values of the ground truth scores and the predicted scores, respectively. KLCC. measures the ordinal association between two measured quantities and is defined as:
where is the number of concordant pairs and is the number of discordant pairs among all possible pairs of data points. A higher KLCC indicates a stronger rank correlation between the two variables. Together, these metrics provide a comprehensive evaluation of the relationship between predicted preference scores and ground truth MOS values across different aspects of correlation.
12 More Details of Our F-Eval Model
12.1 Loss Funtion
We use both language loss, L1 loss and cross-entropy loss as the loss functions to optimize the training process. Specifically, the language loss is used to restrict the F-Eval to produce specific quality-related answer patterns. The language loss function can be formulated as:
where is the predicted token, is the ground truth token, indicates the probability, and is the number of tokens. L1 loss is used to regress the quality scores. The L1 loss can be formulated as:
where is the predicted quality score, is the ground truth quality score, and is the number of images in a batch. Cross-entropy loss to predict the dimension ID from the input text tokens. The cross-entropy loss can be formulated as:
where is the dimension label for the -th sample in class (dimension ID), is the predicted probability for the -th sample in class , and is the number of images in a batch. is 4. The overall loss function can be formulated as:
| Task | Face Generation | Face Restoration syn | |||||||||||||
| Dimension | Quality | Authenticity | Correspondence | Quality | ID Fidelity | ||||||||||
| Method | SRCC | KRCC | PLCC | SRCC | KRCC | PLCC | SRCC | KRCC | PLCC | SRCC | KRCC | PLCC | SRCC | KRCC | PLCC |
| Freeze Projector | 0.8080 | 0.6428 | 0.8698 | 0.7518 | 0.5832 | 0.8123 | 0.8475 | 0.6783 | 0.9031 | 0.7769 | 0.5787 | 0.8074 | 0.7757 | 0.5832 | 0.8074 |
| Freeze Face encoder | 0.8151 | 0.6450 | 0.8733 | 0.7524 | 0.5834 | 0.8140 | 0.8347 | 0.6654 | 0.8912 | 0.8121 | 0.6120 | 0.8430 | 0.7828 | 0.6021 | 0.8341 |
| Single LoRA | 0.7927 | 0.6140 | 0.8659 | 0.7757 | 0.5995 | 0.8285 | 0.8191 | 0.6429 | 0.8760 | 0.8377 | 0.6538 | 0.8874 | 0.8228 | 0.6341 | 0.8556 |
| Task LoRA | 0.8231 | 0.6477 | 0.8895 | 0.7803 | 0.6056 | 0.8317 | 0.8377 | 0.6661 | 0.8943 | 0.8244 | 0.6428 | 0.8806 | 0.8257 | 0.6359 | 0.8519 |
| F-Eval (Ours) | 0.8486 | 0.6670 | 0.9085 | 0.8312 | 0.6585 | 0.8578 | 0.8471 | 0.6637 | 0.9106 | 0.8692 | 0.6855 | 0.9009 | 0.8507 | 0.6731 | 0.8726 |
| Task | Face Customization | Face Restoration rw | |||||||||||||
| Dimension | Quality | Authenticity | Correspondence | Quality | ID Fidelity | ||||||||||
| Method | SRCC | KRCC | PLCC | SRCC | KRCC | PLCC | SRCC | KRCC | PLCC | SRCC | KRCC | PLCC | SRCC | KRCC | PLCC |
| Freeze Projector | 0.9242 | 0.7651 | 0.9233 | 0.9269 | 0.7679 | 0.9242 | 0.8429 | 0.6726 | 0.8386 | 0.7067 | 0.5249 | 0.7114 | 0.4778 | 0.3616 | 0.5451 |
| Freeze Face encoder | 0.9239 | 0.7662 | 0.9217 | 0.9214 | 0.7612 | 0.9208 | 0.8550 | 0.6836 | 0.8524 | 0.7158 | 0.5421 | 0.8532 | 0.6028 | 0.4386 | 0.6898 |
| Single LoRA | 0.9419 | 0.7938 | 0.9421 | 0.9424 | 0.7957 | 0.9430 | 0.8850 | 0.7140 | 0.8974 | 0.8019 | 0.6155 | 0.8455 | 0.6710 | 0.5017 | 0.7768 |
| Task LoRA | 0.9421 | 0.7984 | 0.9403 | 0.9422 | 0.7985 | 0.9402 | 0.8727 | 0.7005 | 0.8918 | 0.7540 | 0.5661 | 0.7964 | 0.6376 | 0.4736 | 0.7403 |
| F-Eval (Ours) | 0.9462 | 0.7961 | 0.9461 | 0.9188 | 0.7640 | 0.9322 | 0.9460 | 0.7959 | 0.9457 | 0.8448 | 0.6577 | 0.8705 | 0.7957 | 0.6057 | 0.8366 |
12.2 Failure Cases
F-Eval’s performance may degrade when the face region in the image is either too small or in an extreme side view. In these scenarios, the face features extracted by the face encoder become less accurate, leading to a drop in performance, particularly in the quality and authenticity dimensions. We will address this limitation by integrating a more robust face encoder fine-tuned on side view and small face data in future works. Additionally, F-Eval currently does not support identifying specific distorted regions, which will also be addressed in future works.
12.3 More Experimental Results
The detailed ablation study results are listed in Tab. 8. The results indicate the effectiveness of the key components in our F-Eval.